### v2

# S2CFormer: Reorienting Learned Image Compression from Spatial Interaction to Channel Aggregation

Abstract
Transformers have achieved significant success in learned image compression (LIC), with Swin Transformers emerging as the mainstream choice for nonlinear transforms. A common belief is that their sophisticated spatial operations contribute most to their efficacy. However, the crucial role of the feed-forward network (FFN) based Channel Aggregation module within the transformer architecture has been largely overlooked, and the over-design of spatial operations leads to a suboptimal trade-off between decoding latency and R-D performance. In this paper, we reevaluate the key factors behind the competence of transformers in LIC. By replacing spatial operations with identity mapping, we are surprised to find that channel operations alone can approach the rate-distortion (R-D) performance of the leading methods. This solid lower bound of performance emphasizes that the presence of channel aggregation is more essential for the LIC model to achieve competitive performance, while the previously complex spatial interactions are partly redundant. This motivates us to use simplified spatial operations and powerful channel operations to overcome the previous trade-off. Based on this insight, we initiate the “S2CFormer” paradigm, a general architecture that reorients the focus of LIC from Spatial Interaction to Channel Aggregation. We present two instantiations of the S2CFormer: S2C-Conv, and S2C-Attention. Each one incorporates a simple operator for spatial interaction and serves as nonlinear transform blocks for our LIC models. Both models demonstrate state-of-the-art (SOTA) R-D performance and significantly faster decoding speed, as shown in Fig. 1. These results also motivate further exploration of advanced FFN structures to enhance the R-D performance while maintaining model efficiency. With these foundations, we introduce S2C-Hybrid, an enhanced LIC model that combines the strengths of different S2CFormer instantiations. This model outperforms all the existing methods on the Kodak, Tecnick, and CLIC Professional Validation datasets, setting a new benchmark for efficient and high-performance LIC. The code is at https://github.com/YunuoChen/S2CFormer.
Figure 1: S2CFormer and the performance of S2CFormer-based models. The general structure of our S2CFormer is shown in (a). It consists of two key components: the Spatial Interaction module and the Channel Aggregation module. S2CFormer functions as nonlinear transform blocks for Learned Image Compression (LIC). We argue that the competence of transformer-based LIC models primarily stems from the Channel Aggregation module, rather than the specific Spatial Interaction module. As illustrated in (b), S2CFormer-based models not only outperform leading LIC models in rate-distortion performance but also offer significant improvements in decoding latency.
1 Introduction
Image compression is crucial due to the rapid growth of digital image data. Lossy compression reduces file sizes while preserving visual quality and optimizing storage and transmission. Recently, learned image compression (LIC) models have emerged as a promising alternative to traditional codecs, offering improved rate-distortion performance and greater efficiency in image storage.
For LIC, nonlinear transform blocks and entropy models are pivotal to a model’s performance. Early LIC models primarily relied on convolutional neural networks (CNNs) for transform networks [10, 8, 41, 3, 20, 40, 49, 38]. However, with the advancements in transformer-based architectures [47, 33], recent studies have adopted transformers as the foundation, leading to significant improvements in rate-distortion (R-D) performance [56, 32, 27, 25, 42, 44, 18, 55].
The success of transformers, particularly Swin Transformers, is often attributed to their sophisticated spatial operations, such as self-attention and window shifting. However, the significance of the transformer structure has not been thoroughly explored, making such attributions premature. Besides, their heavy reliance on spatial operations leads to a suboptimal trade-off between decoding latency and R-D performance.
To this end, we reevaluate the key factors in the R-D performance of transformers-based LIC models. We identify that a basic vision transformer block [47, 33, 34] consists of two operations targeting different dimensions: an attention-based spatial operation and an MLP-based channel operation. By replacing spatial operations with identity mapping, we are surprised to find that channel operations alone can achieve promising performance, which is comparable to the leading methods. The solid lower performance bound and the simple identity spatial operation emphasize that the MLP-based channel operation plays an essential role in optimizing R-D performance for LIC models, while previously complex spatial interactions are partly redundant. This result also gives us hope of overcoming the trade-off between R-D performance and decoding latency inherent in previous methods. To reorient the focus of LIC from Spatial Interaction to Channel Aggregation, we introduce the paradigm “S2CFormer”. S2CFormer is a general structure originated from the standard vision transformers [47, 33, 34], which consists of two main components: Spatial Interaction and Channel Aggregation. This paradigm emphasizes two key ideas: 1) simplifying spatial interactions to accelerate decoding, and 2) aggregating anisotropic features along the channel dimension to achieve promising performance. It realigns the priority of the spatial operator and channel operator to break the limitations of the previous speed-performance trade-off.
Based on the insight of S2CFormer, we adopt two basic operators for spatial interaction: separable convolution, and vanilla attention. These operators, combined with an MLP-based FFN for channel aggregation, constitute two basic S2CFormer instantiations: S2C-Conv, and S2C-Attention. S2C-Conv and S2C-Attention demonstrate that combining simple spatial interaction with channel aggregation can achieve SOTA performance and improve decoding speed by over 30%. Their performances indicate that channel aggregation is the key factor in R-D performance, far outweighing the choice of specific spatial interaction methods (such as convolution or attention). We suppose this is due to the fact that as the resolution of the feature decreases, spatial redundancy shifts to channels, making it crucial to capture and eliminate channel redundancy.
These findings motivate further exploration of advanced FFN structures for channel aggregation, which could enhance LIC models with minimal added complexity. Additionally, utilizing the advantages of different S2C-instantiations, we propose a new LIC model, S2C-Hybrid. This model combines the strengths of both attention and convolution mechanisms for spatial interaction, while employing the advanced Gated-FFN for channel aggregation, thereby significantly improving R-D performance. Besides, it also ensures a much faster decoding speed compared to existing SOTA methods.
Our contributions can be summarized as follows:
  • $\bullet$
    Revealing the Importance of Channel Aggregation in LIC: We emphasize the critical role of channel aggregation in LIC and point out that the previous excessive focus on spatial interactions leads to a suboptimal trade-off between decoding latency and R-D performance.
  • $\bullet$
    Introduction of the S2CFormer Paradigm: We propose S2CFormer to reorient the focus of LIC from spatial interactions to channel aggregation. This paradigm advocates for simplifying spatial interaction and enhancing channel aggregation to improve both speed and performance.
  • $\bullet$
    Establishing New Performance Benchmarks: The S2C-Hybrid model combines the strengths of different S2C-instantiations and outperforms all the existing methods on three datasets (i.e., Kodak, Tecnick, and CLIC datasets), establishing a new R-D performance record for LIC.
2 Related Work
2.1 Learned Image Compression
In the past few years, end-to-end learned image compression (LIC) has gained increased attention. Ballé et al. [2] proposed the first end-to-end image compression model with convolution neural network (CNN) and introduced the variation auto-encoder (VAE) model combined with hyper-prior [3], which has become the most fundamental structure for LIC models. Afterward, researchers mainly focus on two things to improve the rate-distortion performance: transform networks and entropy model [52, 39, 37, 35, 31, 17, 16, 4].
Transform networks refer to the nonlinear analysis and synthesis transforms in Encoder, Decoder, Hyper Encoder, and Hyper Decoder. Cheng et al. [10] adopt Residual blocks. Chen et al. [8] optimize it with octave residual modules. Xie et al. [49] utilized an invertible neural network for better performance. With the development of various transformers, they are also adopted by LIC models. [57, 56] directly utilized Swin Transformer for transforms. TCM [32] employed both Resblock and Swin Transformer to capture both local and global information. FTIC [27] enhances the window attention mechanism from a frequency perspective by carefully designing the attention window.
For the entropy model, researchers mainly focus on exploiting efficient and powerful context models. Minnen et al. [41] proposed an autoregressive entropy model. He et al. [19] utilized the checkerboard model to improve the speed. Minnen et al. [40] proposed a context model along channel dimension. He et al. [20] raised ELIC for unevenly grouped space-channel contextual adaptive coding. [28, 29] proposed a quadtree entropy model for efficiency. Qian et al. [42] utilized ViT for the entropy model for powerful context. MLIC and MLIC++ [22, 21] proposes a multi-reference entropy model. FTIC [27] propose a transformer-based channel-wise autoregressive (T-CA) model that effectively exploits channel dependencies. In this paper, we mainly focus on the structure of transform networks, and directly adopt the basic efficient and powerful entropy model from ELIC [20].
2.2 Transformer
Transformers, initially introduced by [47] for translation tasks, has quickly become widely adopted in a range of NLP applications. Building on this success, many researchers have expanded the use of attention mechanisms and transformers to tackle computer vision tasks. [15, 33, 46, 53, 9].
In particular, Liu et al. [33] introduced the Swin Transformer, a hierarchical model that employs window attention and window shifting mechanism. This innovative structure has established a new performance benchmark in both high-level [48, 14, 34] and low-level [30, 53, 9, 54] vision tasks, reflecting its broad applicability and effectiveness.
Many learned image compression models have adopted the Swin Transformer for nonlinear transform, achieving remarkable improvements in rate-distortion performance. The success of those advanced LIC models has been long attributed to the sophisticated attention modules. However, in the field of natural language processing (NLP), Thorp et al. suggest in their work [26] that substituting the attention module with a simple Fourier transform can yield comparable results. Similarly, in high-level vision tasks, Yu et al. [50, 51] demonstrate that the architecture known as MetaFormer is actually the most crucial factor for achieving high performance.
In this paper, we revisit the competence of transformers in LIC models, arguing that the core factor for superior R-D performance is Channel Aggregation, rather than the well-designed attention modules.
3 Methodology
3.1 Problem Formulation
Figure 2: Overview of S2CFormer-based LIC model. We adopt the basic VAE structure from [41, 3] and integrate the SCCTX entropy model from [20]. The hierarchical architecture consists of five stages of nonlinear transform blocks. Each stage contains $L_{i}$ S2CFormer blocks. The general S2CFormer architecture is shown in (a), and (b-d) illustrate three S2CFormer instantiations. $L_{1}$-$L_{6}$ represent block numbers for each stage.
The structure of our S2CFormer-based LIC model is illustrated in Fig. 2. It consists of three main components: an Encoder, a Decoder, and an Entropy Model. The Encoder generates a latent representation $\bm{y}$ from RGB input $\bm{x}$, which is quantized to $\hat{\bm{y}}$. The Decoder maps $\hat{\bm{y}}$ back to the RGB domain. This process involves an encoder network $g_{a}(\cdot)$, a decoder network $g_{s}(\cdot)$, and a quantization operator $Q$. The steps are:
$\bm{y}=g_{a}(\bm{x};{\bm{\theta}}_{a}),\quad\hat{\bm{y}}=Q(\bm{y}-\bm{\mu})+\bm{\mu},\quad\hat{\bm{x}}=g_{s}(\hat{\bm{y}};{\bm{\theta}}_{s}),$
where $\bm{\mu}$ is the estimated entropy parameter. The entropy model uses a Hyper Encoder $h_{a}(\cdot)$ and a Hyper Decoder $h_{s}(\cdot)$ to transform $\bm{y}$ into a hyper-prior $\bm{z}$, which is quantized and transmitted as side information. The decoder network $h_{s}$ converts $\hat{\bm{z}}$ into gaussian distribution parameters $(\bm{\mu},\bm{\sigma})$:
$\bm{z}=h_{a}(\bm{y};{\bm{\phi}}_{a}),\quad\hat{\bm{z}}=Q(\bm{z}),\quad\bm{\mu},\bm{\sigma}=h_{s}(\hat{\bm{z}};{\bm{\phi}}_{s}).$
The model is trained to minimize the R-D loss function:
$\mathcal{L}=$ $\mathcal{R}(\hat{\bm{y}})+\mathcal{R}(\hat{\bm{z}})+\lambda\cdot\mathcal{D}(\bm{x},\hat{\bm{x}})$
$=$ $\mathbb{E}\left[-\log_{2}p_{\hat{\bm{y}}\mid\hat{\bm{z}}}(\hat{\bm{y}}\mid\hat{\bm{z}})\right]+\mathbb{E}\left[-\log_{2}p_{\widehat{\bm{z}}}(\widehat{\bm{z}})\right]$
$+\lambda\cdot\mathbb{E}\|\bm{x}-\widehat{\bm{x}}\|_{2}^{2},$
where $\mathcal{R}(\hat{\bm{y}})$ and $\mathcal{R}(\hat{\bm{z}})$ are bitrates estimated by the entopy model, and $\mathcal{D}(\bm{x},\hat{\bm{x}})$ is the distortion between the original and reconstructed images, with $\lambda$ balancing bitrate and distortion.
We focus on the design of four transform networks: Encoder $g_{a}$, Decoder $g_{s}$, Hyper Encoder $h_{a}$, and Hyper Decoder $h_{s}$, each including several nonlinear transform blocks. The S2CFormer block is employed for these transforms. Downsampling is achieved with a $3\times 3$ convolution layer (stride 2), and upsampling uses a pixel shuffle layer with a $3\times 3$ convolution. The entropy models adopt the efficient Space-Channel ConTeXt (SCCTX) model from ELIC [20], enhancing entropy estimation with spatial and channel context.
3.2 S2CFormer
3.2.1 S2C-Identity: The Lower Bound of S2CFormers
Firstly, we introduce the “S2CFormer” paradigm, shown in Fig. 2 (a). S2CFormer is a general architecture comprising two main components: Spatial Interaction and Channel Aggregation.
To reevaluate the key factors in the R-D performance of transformers-based LIC models and explore the lower bound of S2CFormer’s performance, we instantiate the spatial interaction with the simplest operator, Identity Mapping, as illustrated in Fig. 2 (b):
$\operatorname{Identity}({\bm{X}})={\bm{X}},$
which is an identity spatial transformation, and does not perform special spatial feature interaction. We use this initialization, termed S2C-Identity, to demonstrate the effectiveness and dominance of channel aggregation.
3.2.2 Simplified Spatial Interaction
In addition to S2C-Identity, we explore more S2CFormer instantiations with other spatial operators, as shown in Fig. 2 (c-d). Instead of designing novel spatial interaction modules, we assess the model’s potential for achieving SOTA performance with simplified, commonly used spatial operators. We implement two spatial interaction modules. The first is separable convolution, as proposed in [50, 11, 45]:
$\operatorname{SepConv}({\bm{X}})=\operatorname{Conv}_{\mathrm{pw}}\left(\operatorname{Conv}_{\mathrm{dw}}\left(\sigma\left(\operatorname{Conv}_{\mathrm{pw}}({\bm{X}})\right)\right)\right),$
where $\operatorname{Conv}_{\mathrm{pw}}$ and $\operatorname{Conv}_{\mathrm{dw}}$ denote point-wise and depth-wise convolution, respectively, and $\sigma$ represents nonlinear activation function. Separable convolutions help to reduce computational cost and parameter count. This instantiation is termed as S2C-Conv.
Besides convolution, another commonly used spatial interaction module for LIC is self-attention. We employ a simplified one, vanilla window attention, to achieve basic spatial interactions. It is defined as:
$\operatorname{Attention}({\bm{Q}},{\bm{K}},{\bm{V}})=\operatorname{Softmax}\left(\frac{{\bm{Q}}{\bm{K}}^{\top}}{\sqrt{d_{h}}}\right){\bm{V}},$
where ${\bm{Q}},{\bm{K}},{\bm{V}}$ represent the query, key, and value, respectively, and $d_{h}$ denotes the number of attention heads. It computes attention in a windowed manner, eliminating special window partitions or time-consuming window shifting operations. This instantiation is identified as S2C-Attention.
Following spatial interaction, we apply channel aggregation with Layer Normalization and a feed-forward network (FFN) consisting of linear layers, as detailed in Sec. 3.2.3.
3.2.3 Advanced Channel Aggregation
This paper does not introduce complex spatial operations. Instead, we explore advanced FFN structures for channel aggregation to assess their potential benefits, aiming to enhance LIC models with minimal added complexity and achieve a better balance between performance and coding speed. Channel aggregation is primarily achieved using MLP-based FFNs. As illustrated in Fig. 3 (a), a vanilla FFN can be formulated as follows:
$\operatorname{Vanilla\text{-}FFN}({\bm{X}})=\sigma\left({\bm{X}}{\bm{W}}_{\rm in}\right){\bm{W}}_{\rm out},$
where ${\bm{W}}_{\rm in}\in\mathbb{R}^{C\times rC}$ and ${\bm{W}}_{\rm out}\in\mathbb{R}^{rC\times C}$ are learnable parameters with an expansion ratio $r$ and the input channel number $C$. $\sigma$ refers to the activation function.
Inspired by previous work in NLP and low-level tasks [9, 53, 13, 29, 28], we propose two efficient yet powerful FFN structures for channel aggregation: Additive-FFN and Gated-FFN, shown in Fig. 3 (b) and (c).
Figure 3: Vanilla FFN (a) and Advanced FFNs (b-c)
The Additive-FFN combines outputs from two activation functions via addition, applying separate nonlinear transformations to the same input to leverage the strength of multiple nonlinearities:
$\operatorname{Additive\text{-}FFN}({\bm{X}})=\left[\sigma_{1}\left({\bm{X}}{\bm{W}}_{1}\right)+\sigma_{2}\left({\bm{X}}{\bm{W}}_{2}\right)\right]{\bm{W}}_{\text{out}},$
where ${\bm{W}}_{1}$ and ${\bm{W}}_{2}$ are learnable parameters, and $\sigma_{1}$, $\sigma_{2}$ are distinct activation functions.
In contrast, the Gated-FFN enhances feature interactions via the Hadamard product, performing element-wise multiplication of two linear transformation results, with one passed through an activation function:
$\operatorname{Gated\text{-}FFN}({\bm{X}})=\left[\sigma\left({\bm{X}}{\bm{W}}_{1}\right)\odot\left({\bm{X}}{\bm{W}}_{2}\right)\right]{\bm{W}}_{\text{out}},$
where $\odot$ denotes element-wise multiplication. For all structures, the expansion ratio $r$ is set to 4, and linear layers are implemented via $1\times 1$ convolutions.
These advanced FFNs outperform the vanilla FFN while maintaining similar computational complexity and latency. This suggests that exploring efficient FFN designs is crucial for effective LIC models, beyond relying solely on complex spatial interactions.
3.2.4 S2C-Hybrid: A Superior Extension of S2CFormers
Building on the spatial interaction operators described in Sec. 3.2.2 and the advanced channel aggregation techniques from Sec. 3.2.3, we introduce the S2C-Hybrid model, an enhanced version of the S2CFormer-based LIC model. S2C-Hybrid offers an excellent trade-off between R-D performance and decoding speed.
This hybrid model employs different S2C instantiations for different transform stages. For channel aggregation, we adopt gated-FFN to enhance performance. As for spatial interaction, S2C-Hybrid takes advantage of both convolutional networks and self-attention mechanisms. Specifically, we employ convolution in the early stages when feature resolution is high, and apply attention mechanisms in the later stages. This strategy improves R-D performance without significantly increasing time complexity or slowing down coding speed. Additionally, the S2C-Conv and S2C-Identity blocks are used to enhance spatial context extraction and parameter aggregation in the SCCTX entropy model.
With advanced channel aggregation to enhance performance and simplified spatial interactions to accelerate speed, S2C-Hybrid overcomes the suboptimal trade-off between performance and speed inherent in previous methods.
ELIC MambaVC FTIC MLIC++ TCM-large S2C-Identity S2C-Conv S2C-Attention S2C-Hybrid
BD-rate [Kodak] -3.10 -8.11 -12.94 -11.97 -10.04 -9.70 -12.65 -12.56 -14.28
BD-rate [CLIC] -0.84 - -10.21 -12.08 -8.60 -7.55 -10.98 -10.28 -12.88
BD-rate [Tecnick] -7.41 - -13.89 -15.13 -10.42 -10.70 -14.48 -14.08 -17.20
Params (M) 33.29 47.88 69.78 116.48 75.89 64.63 66.60 68.42 79.83
FLOPs (T) 1.74 2.10 2.38 2.64 3.74 2.80 3.13 3.42 4.17
[…]
Throughput (samples/s) $\uparrow$ 72.32 6.55 23.25 27.42 17.80 50.02 35.49 26.21 22.63
Table 1: Comprehensive comparisons of LIC models. (a) BD-rate. BD-rate of LIC models on different datasets, relative to VTM-21.0 [7]. S2C-Conv, -Attention, and -Hybrid models achieve superior R-D performance. R-D curves are provided in Fig. 6-8. (b) Model Params. S2C-Conv and -Attention models have fewer parameters compared to previous SOTA methods. (c) Efficiency. FLOPs assess computational complexity, but they do not always reflect the actual training and inference speed. While S2CFormer-based models do not exhibit many advantages in terms of FLOPs, they achieve over 30% faster decoding speed and significantly higher training throughput.
4 Experiments
4.1 Experimental Setup
4.1.1 Training Details
We train all S2CFormer-based models on the Flickr2W dataset [31] for 2 million steps with a batch size of 8, using Adam [23] optimizer and an initial learning rate of 0.0001. For MSE-optimized models, we use Lagrangian multipliers {0.0017, 0.0025, 0.0035, 0.0067, 0.0130, 0.0250, 0.050}; for MS-SSIM-optimized models, we use {3, 5, 8, 16, 36, 64}. In the attention modules, the window size is set to 8, and depthwise convolution kernels are all of size 5. The latent representation $y$ consists of 320 channels, and the hyper-latent representation contains 192 channels. Additional details are provided in the supplementary material. To enhance context mining and parameter aggregation within the entropy model of ELIC [20], we incorporate the S2C-Conv block and the S2C-Identity block, respectively. Unless otherwise specified, we use the Gated-FFN as a channel aggregation module for all S2CFormer-based models. All experiments are conducted on NVIDIA A100 GPUs.
4.1.2 Evaluation
We test our models on three datasets: Kodak image set [24] with the image size of 768 × 512, Tecnick test set [1] with the image size of 1200 × 1200, and CLIC professional validation dataset [12] with 2k resolution. We take PSNR, MS-SSIM, and bits per pixel (bpp) as metrics. We utilize the BD-rate [5] results to quantify the average bitrate savings.
4.2 Rate-Distortion Performance
In this section, we compare our models with VTM-21.0 [7, 6] and several leading LIC models, including ELIC [20], MambaVC [43], TCM [32], MLIC++ [22] and FTIC [27]. The detailed results are presented in Table 1.
The Lower Bound of S2CFormers.  S2C-Identity, which relies mainly on MLPs without spatial interactions, holds the lower performance bound for S2CFormer-based models. Surprisingly, it outperforms VTM-21.0 by 9.70%, 10.70%, and 7.55% in BD-rate on Kodak, Tecnick, and CLIC datasets, respectively. It surpasses MambaVC, and nearly matches the performance of TCM-large, as shown in Fig. 4. This promising result suggests that MLP-based channel aggregation secures a solid baseline performance for LIC.
S2CFormers With Simplified Spatial Interaction.  We compare the S2C-Conv and S2C-Attention models with existing models and find that both models outperform nearly all the SOTA methods. As shown in Tabel 1, in terms of BD-rates, both models surpass FTIC, TCM-large, and MambaVC. They also outperform MLIC++ on the Kodak dataset. Although MLIC++ performs well at lower bit rates, it underperforms at higher bitrates, showing a noticeable gap compared to our models, as shown in Fig. 4. This discrepancy arises from its focus on improving the entropy model, which benefits performance at lower bitrates. However, at higher bit rates, transform networks become more critical for capturing and reconstructing high-frequency components.
The performances of our models suggest that complex spatial interaction design may be redundant and can be simplified. S2C-Conv and S2C-Attention show near-identical performance, indicating a specific choice of spatial interaction is not critical for LIC. Instead, FFN-based channel aggregation is the key factor that secures R-D performance.
Figure 4: Ablation Study for Channel Aggregation. Experiments on Kodak dataset. “w/ FFN” refers to S2CFormer-based models with Channel Aggregation. “w/o CA” represents removing all FFN modules for Channel Aggregation.
Figure 5: The effective receptive fields (ERF) [36] calculated by different models. “CA” refers to Channel Aggregation module.
Figure 6: Performance evaluation on the CLIC Professional Validation Dataset.
Figure 7: Performance evaluation on the Kodak dataset.
Figure 8: Performance evaluation on the Tecnick dataset.
Ablations of Channel Aggregation.  In this section, we further analyze the critical role of Channel Aggregation within our S2CFormer architecture. To assess their impact, we first conducted an ablation study by removing all FFN modules from S2C-Conv and S2C-Attention.
Fig. 4 illustrates the R-D curves for this ablation study. Notably, pure separable convolution and window attention are significantly less effective than their FFN-enhanced counterparts. Specifically, adding FFNs to separable convolution improves BD-rate by -10.57%, while FFNs in S2C-Attention lead to an -11.14% improvement. These results highlight the substantial contribution of FFNs to the overall R-D performance of our LIC models.
In contrast, S2C-Identity, which eliminates spatial interaction while retaining channel aggregation, approaches the performance of TCM-large, outperforming the two models without CA by -7.51% and -8.17% BD-rate respectively. This comparison underscores the greater importance of FFNs in LIC models.
This ablation study strongly supports using FFNs for channel aggregation in future LIC models.
Figure 9: Comparison of Different FFNs. Experiments on the Kodak dataset. We provide R-D curves for S2C-Conv and -Attention models with various FFNs for Channel Aggregation.
Figure 10: Visual Comparison. This figure presents visualizations of decompressed images from the Kodak dataset using various methods. Each subfigure is labeled with “Method |||| Bpp |||| PSNR |||| MS-SSIM”.
We also examine the ERF [36] of analysis networks to demonstrate the significance of channel aggregation. Although it is commonly believed that the attention mechanism expands the receptive field and aggregates more non-local information, our findings challenge this view. As shown in Fig. 5, both the S2C-Conv and S2C-Attention models, when lacking channel aggregation, exhibit smaller ERF. In contrast, when channel aggregation is introduced, the ERF of both models expands to cover more areas of the input image. We also evaluate the Swin Transformer-based LIC model and observe that window shifting does not significantly enlarge the ERF. Similar to the S2C-based models, Swin-based models also rely on channel aggregation to broaden ERF. This demonstrates that channel aggregation is the key factor in activating more pixels in the original image and capturing additional information, which boosts the R-D performance of LIC models. These comparisons suggest that some complex and time-consuming spatial operations appear redundant.
Effectiveness of Advanced Channel Aggregation.  As discussed in Sec. 3.2.3, the structure of FFN warrants further exploration. We conduct comparison experiments on three proposed FFN structures—Vanilla-FFN, Additive-FFN, and Gated-FFN—within both the S2C-Conv and S2C-Attention models, as shown in Fig. 9. For S2C-Conv, models utilizing Vanilla-FFN, Additive-FFN, and Gated-FFN outperform VTM-21.0 by -10.88%, -11.87%, and -12.65%, respectively. When window attention is applied for spatial interaction, these models achieved BD-rates of -11.27%, -12.05%, and -12.56% over VTM-21.0. The decoding latency for these models is nearly identical. Gated-FFN consistently outperforms Additive-FFN, which, in turn, surpasses Vanilla-FFN. These results demonstrate that the FFN structure remains a promising area for further investigation.
Performance of S2C-Hybrid.  We enhance model performance by integrating the advantages of different S2C-instantiations and advanced FFN. We increase $L_{2}$ and $L_{3}$ to 8 while keeping other configurations unchanged. To manage decoding latency and complexity, we utilize the S2C-Conv block in the initial stage to reduce complexity; in subsequent stages, we employ the S2C-Attention block. As shown in Fig. 6-8, by employing only the fundamental S2CFormer blocks with simplified spatial interaction modules, our S2C-Hybrid model surpasses all other LIC models in R-D performance, exceeding VTM-21.0 by 14.28%, 12.88%, and 17.20% in BD-rate for the Kodak, CLIC, and Tecnick datasets, respectively. This result demonstrates the effectiveness of integrating various simple spatial interactions within a unified architecture to optimize both computational efficiency and model performance.
Ablations on S2C-Hybrid For the three-stage main transform networks, we employ S2C-Conv for the first stage and S2C-Attention for the others, which is the best arrangement according to our experiments. As shown in Tab.2, we represent S2C-Conv and -Attention with “C” and “A” respectively. “[C,A,A]” appears to be the best, the reason for which may be that convolution can better capture local details at high resolution, while attention excels at global information when resolution drops.
Stages[1,2,3] [C,C,C] [A,A,A] [A,C,C] [C,C,A] [C,A,A]
BD-rate (%) -12.65 -12.56 -12.32 -12.87 -13.35
Table 2: S2C-Hybrid Arrangement Ablations
We also have experimented about the S2C-Hybrid model’s complexity in Tab.3. The base model Hybrid-S has 3 blocks for each stage. By stacking more blocks in the last two stages (adding blocks to the first stage greatly slows decoding), performance steadily improves.
Stages[1,2,3] Hybrid-S[3,3,3] Hybrid-M[3,5,5] Hybrid-L[3,8,8]
BD-rate (%) -13.35 -13.85 -14.28
FLOPs (T) 3.20 3.59 4.17
Decoding (s) $\downarrow$ 0.357 0.358 0.362
Table 3: S2C-Hybrid Complexity Ablations
Visual Comparison.  We present the visualization results in Fig. 10, comparing our method with the traditional codec VTM 21.0 and the advanced LIC model TCM.
4.3 Model efficiency
We calculate the FLOPs and decoding latency for various models using 2K images, and evaluate training throughput with a batch size of 8 and patch size of 256×256. Training throughput is directly correlated to the training time of LIC models. Numeric results are provided in Table 1.
Decoding Latency.  S2CFormer-based models demonstrate significant advantages in decoding latency. All of our models are quite fast, with even the slowest one, S2C-Hybrid, achieving a decoding speed that is more than 30% faster than MLIC++ and TCM-large.
Training Throughput.  S2CFormer-based models also significantly increase training throughput, S2C-Identity is competitive with TCM-large in R-D performance, while improving training speed by 180%, which cuts the total training time from 10.4 days to 3.7 days. Similarly, S2C-Conv outperforms FTIC in R-D performance, and accelerates training by 52%, saving over 65 hours of training time.
FLOPs.  For VAE-based LIC models, FLOPs are significantly influenced by the number of channels in the first stage. We directly set all channel numbers as 192, following the original setting of LIC model [2], as the optimal channel settings are beyond the scope of this study. Some leading methods have reduced the first-stage channel number to 128 or fewer [43, 27], which significantly lowers FLOPs. However, these models are still much slower than our S2CFormer-based models in both training and decoding speed, which are the primary concerns for LIC.
5 Conclusion and Future Work
We revisit the key factors that enhance the R-D performance of transformer-based LIC models. Specifically, we identify certain spatial operations in previous leading methods as redundant. To address this, we introduce S2CFormer, a novel paradigm for nonlinear transform blocks in LIC. This architecture consists of two core components: the Spatial Interaction module and the Channel Aggregation module. By shifting the focus of LIC from spatial interaction to channel aggregation, this approach aims to optimize performance while also improving speed. Through extensive experiments and comparisons, we demonstrate that FFN-based channel aggregation is critical for improving LIC performance. Models based on S2CFormer offer faster training speed and lower decoding latencies compared to existing methods while maintaining competitive R-D performance. Moreover, we suggest that FFN-based channel aggregation remains a promising avenue for future research.
By combining simple spatial interaction operators with the proposed Gated-FFN for channel aggregation, our models achieve state-of-the-art R-D performance on the Kodak, Tecnick, and CLIC Professional Validation datasets, while also maintaining efficiency. This approach breaks the suboptimal trade-off between R-D performance and decoding latency inherent in previous methods.

### v3

@@
-S2CFormer: Reorienting Learned Image Compression from Spatial Interaction to Channel Aggregation
+S2CFormer: Revisiting the RD-Latency Trade-off in Transformer-based Learned Image Compression
 Abstract
-Transformers have achieved significant success in learned image compression (LIC), with Swin Transformers emerging as the mainstream choice for nonlinear transforms.
-A common belief is that their sophisticated spatial operations contribute most to their efficacy.
-However, the crucial role of the feed-forward network (FFN) based Channel Aggregation module within the transformer architecture has been largely overlooked, and the over-design of spatial operations leads to a suboptimal trade-off between decoding latency and R-D performance.
-In this paper, we reevaluate the key factors behind the competence of transformers in LIC.
-By replacing spatial operations with identity mapping, we are surprised to find that channel operations alone can approach the rate-distortion (R-D) performance of the leading methods.
-This solid lower bound of performance emphasizes that the presence of channel aggregation is more essential for the LIC model to achieve competitive performance, while the previously complex spatial interactions are partly redundant.
-This motivates us to use simplified spatial operations and powerful channel operations to overcome the previous trade-off.
-Based on this insight, we initiate the “S2CFormer” paradigm, a general architecture that reorients the focus of LIC from Spatial Interaction to Channel Aggregation.
-We present two instantiations of the S2CFormer: S2C-Conv, and S2C-Attention.
-Each one incorporates a simple operator for spatial interaction and serves as nonlinear transform blocks for our LIC models.
+Transformer-based Learned Image Compression (LIC) suffers from a suboptimal trade-off between decoding latency and rate-distortion (R-D) performance.
+Moreover, the critical role of the FeedForward Network (FFN)-based channel aggregation module has been largely overlooked.
+Our research reveals that efficient channel aggregation—rather than complex and time-consuming spatial operations—is the key to achieving competitive LIC models.
+Based on this insight, we initiate the “S2CFormer” paradigm, a general architecture that simplifies spatial operations and enhances channel operations to overcome the previous trade-off.
+We present two instances of the S2CFormer: S2C-Conv, and S2C-Attention.
 Both models demonstrate state-of-the-art (SOTA) R-D performance and significantly faster decoding speed, as shown in Fig.
 1.
-These results also motivate further exploration of advanced FFN structures to enhance the R-D performance while maintaining model efficiency.
-With these foundations, we introduce S2C-Hybrid, an enhanced LIC model that combines the strengths of different S2CFormer instantiations.
+Furthermore, we introduce S2C-Hybrid, an enhanced variant that maximizes the strengths of different S2CFormer instances to achieve a better performance-latency trade-off.
 This model outperforms all the existing methods on the Kodak, Tecnick, and CLIC Professional Validation datasets, setting a new benchmark for efficient and high-performance LIC.
@@
 S2CFormer functions as nonlinear transform blocks for Learned Image Compression (LIC).
-We argue that the competence of transformer-based LIC models primarily stems from the Channel Aggregation module, rather than the specific Spatial Interaction module.
-As illustrated in (b), S2CFormer-based models not only outperform leading LIC models in rate-distortion performance but also offer significant improvements in decoding latency.
+Our analysis reveals that the competence of transformer-based LIC models primarily stems from channel aggregation.
+Building on this insight, we propose a novel design strategy that rebalances these two modules to achieve a more favorable trade-off between compression performance and decoding latency.
+As illustrated in (b), the data points for S2CFormer-based models exhibit a linear trend with a steeper slope, thereby underscoring their superior performance–latency characteristics.
 1 Introduction
+Figure 2: Comparison of execution times for spatial interaction and channel aggregation across different models.
+Previous methods show much higher spatial interaction times than channel aggregation, causing significant delays.
+Our S2CFormer effectively rebalances the time relationship between these two modules.
 Image compression is crucial due to the rapid growth of digital image data.
@@
 However, with the advancements in transformer-based architectures [48, 34], recent studies have adopted transformers as the foundation, leading to significant improvements in rate-distortion (R-D) performance [57, 33, 28, 26, 43, 45, 19, 56].
-The success of transformers, particularly Swin Transformers, is often attributed to their sophisticated spatial operations, such as self-attention and window shifting.
+The success of transformers is often attributed to their sophisticated spatial operations, such as self-attention and window shifting.
 However, the significance of the transformer structure has not been thoroughly explored, making such attributions premature.
-Besides, their heavy reliance on spatial operations leads to a suboptimal trade-off between decoding latency and R-D performance.
+Besides, as shown in Fig.
+1 and Fig.
+2, their complex and heavy spatial operations are quite time-consuming, thus often leading to a suboptimal trade-off between decoding latency and R-D performance.
 To this end, we reevaluate the key factors in the R-D performance of transformers-based LIC models.
 We identify that a basic vision transformer block [48, 34, 35] consists of two operations targeting different dimensions: an attention-based spatial operation and an MLP-based channel operation.
-By replacing spatial operations with identity mapping, we are surprised to find that channel operations alone can achieve promising performance, which is comparable to the leading methods.
-The solid lower performance bound and the simple identity spatial operation emphasize that the MLP-based channel operation plays an essential role in optimizing R-D performance for LIC models, while previously complex spatial interactions are partly redundant.
-This result also gives us hope of overcoming the trade-off between R-D performance and decoding latency inherent in previous methods.
+By replacing spatial operations with identity mapping, we are surprised to find that channel operations alone can achieve promising performance, which is even comparable to the leading methods.
+The solid lower performance bound and the simple identity spatial operation emphasize that the MLP-based channel operation plays an essential role in optimizing R-D performance for LIC models, while previously complex spatial interactions may be partly redundant.
 To reorient the focus of LIC from Spatial Interaction to Channel Aggregation, we introduce the paradigm “S2CFormer”.
-S2CFormer is a general structure originated from the standard vision transformers [47, 33, 34], which consists of two main components: Spatial Interaction and Channel Aggregation.
+S2CFormer is a general structure originated from the standard vision transformers, which consists of two main components: Spatial Interaction and Channel Aggregation.
 This paradigm emphasizes two key ideas: 1) simplifying spatial interactions to accelerate decoding, and 2) aggregating anisotropic features along the channel dimension to achieve promising performance.
@@
 Based on the insight of S2CFormer, we adopt two basic operators for spatial interaction: separable convolution, and vanilla attention.
-These operators, combined with an MLP-based FFN for channel aggregation, constitute two basic S2CFormer instantiations: S2C-Conv, and S2C-Attention.
+These operators, combined with a FFN for channel aggregation, constitute two basic S2CFormer instances: S2C-Conv, and S2C-Attention.
 S2C-Conv and S2C-Attention demonstrate that combining simple spatial interaction with channel aggregation can achieve SOTA performance and improve decoding speed by over 30%.
-Their performances indicate that channel aggregation is the key factor in R-D performance, far outweighing the choice of specific spatial interaction methods (such as convolution or attention).
-We suppose this is due to the fact that as the resolution of the feature decreases, spatial redundancy shifts to channels, making it crucial to capture and eliminate channel redundancy.
-These findings motivate further exploration of advanced FFN structures for channel aggregation, which could enhance LIC models with minimal added complexity.
-Additionally, utilizing the advantages of different S2C-instantiations, we propose a new LIC model, S2C-Hybrid.
-This model combines the strengths of both attention and convolution mechanisms for spatial interaction, while employing the advanced Gated-FFN for channel aggregation, thereby significantly improving R-D performance.
+Their performances indicate that channel aggregation is the key factor in R-D performance, far outweighing the choice of specific spatial interaction methods (e.g., convolution or attention).
+We attribute this to reduced feature resolution shifting spatial redundancy to channels, highlighting the importance of capturing and eliminating channel redundancy.
+These findings motivate further exploration of advanced MLP structures for channel aggregation, which could enhance LIC models with minimal added complexity.
+Additionally, we propose a new LIC model, S2C-Hybrid.
+This model achieves even better compression performance by appropriately arranging and combining different S2C-instances, thus fully exploiting the strengths of each S2C-instantiation.
 Besides, it also ensures a much faster decoding speed compared to existing SOTA methods.
@@
   • $\bullet$
-    Establishing New Performance Benchmarks: The S2C-Hybrid model combines the strengths of different S2C-instantiations and outperforms all the existing methods on three datasets (i.e., Kodak, Tecnick, and CLIC datasets), establishing a new R-D performance record for LIC.
+    Establishing New Performance Benchmarks: The S2C-Hybrid model combines the strengths of different S2C-instances and outperforms all the existing methods on three datasets (i.e., Kodak, Tecnick, and CLIC datasets), establishing a new R-D performance record for LIC.
 2 Related Work
@@
 Ballé et al.
-[2] proposed the first end-to-end image compression model with convolution neural network (CNN) and introduced the variation auto-encoder (VAE) model combined with hyper-prior [3], which has become the most fundamental structure for LIC models.
-Afterward, researchers mainly focus on two things to improve the rate-distortion performance: transform networks and entropy model [52, 39, 37, 35, 31, 17, 16, 4].
+[2] proposed the first end-to-end image compression model with convolution neural network (CNN) and introduced the variation auto-encoder (VAE) model combined with hyper-prior [3], which has become the fundamental structure for LIC models.
+Afterward, researchers mainly focus on two things to improve the rate-distortion performance: transform networks and entropy model [53, 40, 38, 36, 32, 17, 16, 4, 18].
 Transform networks refer to the nonlinear analysis and synthesis transforms in Encoder, Decoder, Hyper Encoder, and Hyper Decoder.
@@
 In particular, Liu et al.
-[33] introduced the Swin Transformer, a hierarchical model that employs window attention and window shifting mechanism.
-This innovative structure has established a new performance benchmark in both high-level [48, 14, 34] and low-level [30, 53, 9, 54] vision tasks, reflecting its broad applicability and effectiveness.
-Many learned image compression models have adopted the Swin Transformer for nonlinear transform, achieving remarkable improvements in rate-distortion performance.
-The success of those advanced LIC models has been long attributed to the sophisticated attention modules.
+[34] introduced the Swin Transformer, a hierarchical model that employs window attention and window shifting mechanisms.
+This innovative structure has established a new performance benchmark in both high-level [49, 14, 35] and low-level [31, 54, 9, 55] vision tasks.
+Many LIC models have adopted the Swin Transformer for nonlinear transform, achieving remarkable improvements in RD performance.
+The success of those models has been long attributed to the sophisticated attention modules.
 However, in the field of natural language processing (NLP), Thorp et al.
@@
 Each stage contains $L_{i}$ S2CFormer blocks.
-The general S2CFormer architecture is shown in (a), and (b-d) illustrate three S2CFormer instantiations.
-$L_{1}$-$L_{6}$ represent block numbers for each stage.
+The general S2CFormer architecture is shown in (a), and (b-d) illustrate three S2CFormer instances.
+$L_{1}$-$L_{6}$ and $C_{1}$-$C_{6}$ represent block numbers and channel numbers for each stage, respectively
 The structure of our S2CFormer-based LIC model is illustrated in Fig.
-2.
+3.
 It consists of three main components: an Encoder, a Decoder, and an Entropy Model.
@@
 The Decoder maps $\hat{\bm{y}}$ back to the RGB domain.
-This process involves an encoder network $g_{a}(\cdot)$, a decoder network $g_{s}(\cdot)$, and a quantization operator $Q$.
-The steps are:
+This process involves an encoder network $g_{a}(\cdot)$, a decoder network $g_{s}(\cdot)$, and a quantization operator $Q$:
 $\bm{y}=g_{a}(\bm{x};{\bm{\theta}}_{a}),\quad\hat{\bm{y}}=Q(\bm{y}-\bm{\mu})+\bm{\mu},\quad\hat{\bm{x}}=g_{s}(\hat{\bm{y}};{\bm{\theta}}_{s}),$
@@
 where $\mathcal{R}(\hat{\bm{y}})$ and $\mathcal{R}(\hat{\bm{z}})$ are bitrates estimated by the entopy model, and $\mathcal{D}(\bm{x},\hat{\bm{x}})$ is the distortion between the original and reconstructed images, with $\lambda$ balancing bitrate and distortion.
-We focus on the design of four transform networks: Encoder $g_{a}$, Decoder $g_{s}$, Hyper Encoder $h_{a}$, and Hyper Decoder $h_{s}$, each including several nonlinear transform blocks.
-The S2CFormer block is employed for these transforms.
-Downsampling is achieved with a $3\times 3$ convolution layer (stride 2), and upsampling uses a pixel shuffle layer with a $3\times 3$ convolution.
-The entropy models adopt the efficient Space-Channel ConTeXt (SCCTX) model from ELIC [20], enhancing entropy estimation with spatial and channel context.
 3.2 S2CFormer
+In this paper, we focus on overcoming the previously sub-optimal RD-latency trade-off by rebalancing spatial interaction and channel aggregation.
+Existing studies have typically attributed the success of transformers in RD performance to attention-based spatial interactions, while research on channel aggregation modules for LIC remains limited.
+However, as illustrated in Fig.
+2, excessively complex spatial interaction modules significantly degrade decoding speed, whereas channel aggregation appears comparatively more efficient.
+Consequently, we aim to determine whether channel aggregation might be the primary factor affecting RD performance.
+To this end, we propose a new paradigm called “S2CFormer,” which reorients LIC from spatial interaction towards channel aggregation by significantly simplifying spatial interaction modules.
+As depicted in Fig.
+3 (a), our proposed paradigm is a general architecture consisting of two primary components: Spatial Interaction and Channel Aggregation.
 3.2.1 S2C-Identity: The Lower Bound of S2CFormers
-Firstly, we introduce the “S2CFormer” paradigm, shown in Fig. 2 (a). S2CFormer is a general architecture comprising two main components: Spatial Interaction and Channel Aggregation.
 To reevaluate the key factors in the R-D performance of transformers-based LIC models and explore the lower bound of S2CFormer’s performance, we instantiate the spatial interaction with the simplest operator, Identity Mapping, as illustrated in Fig.
@@
 3.2.2 Simplified Spatial Interaction
-In addition to S2C-Identity, we explore more S2CFormer instantiations with other spatial operators, as shown in Fig.
+In addition to S2C-Identity, we explore more S2CFormer instances with other spatial operators, as shown in Fig.
 3 (c-d).
@@
 where $\odot$ denotes element-wise multiplication. For all structures, the expansion ratio $r$ is set to 4, and linear layers are implemented via $1\times 1$ convolutions.
-These advanced FFNs outperform the vanilla FFN while maintaining similar computational complexity and latency.
-This suggests that exploring efficient FFN designs is crucial for effective LIC models, beyond relying solely on complex spatial interactions.
-3.2.4 S2C-Hybrid: A Superior Extension of S2CFormers
-Building on the spatial interaction operators described in Sec.
-3.2.2 and the advanced channel aggregation techniques from Sec.
-3.2.3, we introduce the S2C-Hybrid model, an enhanced version of the S2CFormer-based LIC model.
-S2C-Hybrid offers an excellent trade-off between R-D performance and decoding speed.
-This hybrid model employs different S2C instantiations for different transform stages.
-For channel aggregation, we adopt gated-FFN to enhance performance.
-As for spatial interaction, S2C-Hybrid takes advantage of both convolutional networks and self-attention mechanisms.
-Specifically, we employ convolution in the early stages when feature resolution is high, and apply attention mechanisms in the later stages.
-This strategy improves R-D performance without significantly increasing time complexity or slowing down coding speed.
-Additionally, the S2C-Conv and S2C-Identity blocks are used to enhance spatial context extraction and parameter aggregation in the SCCTX entropy model.
-With advanced channel aggregation to enhance performance and simplified spatial interactions to accelerate speed, S2C-Hybrid overcomes the suboptimal trade-off between performance and speed inherent in previous methods.
-ELIC MambaVC FTIC MLIC++ TCM-large S2C-Identity S2C-Conv S2C-Attention S2C-Hybrid
-BD-rate [Kodak] -3.10 -8.11 -12.94 -11.97 -10.04 -9.70 -12.65 -12.56 -14.28
-BD-rate [CLIC] -0.84 - -10.21 -12.08 -8.60 -7.55 -10.98 -10.28 -12.88
-BD-rate [Tecnick] -7.41 - -13.89 -15.13 -10.42 -10.70 -14.48 -14.08 -17.20
-Params (M) 33.29 47.88 69.78 116.48 75.89 64.63 66.60 68.42 79.83
-FLOPs (T) 1.74 2.10 2.38 2.64 3.74 2.80 3.13 3.42 4.17
+Advanced FFNs surpass vanilla FFNs in performance with comparable computational complexity and latency, highlighting the importance of efficient FFN designs for effective LIC models, beyond complex spatial interactions.
+3.2.4 S2C-Hybrid: A Superior Extension.
+The strong RD performance of S2C-Conv and S2C-Attention highlights the redundancy of prior complex spatial operations but raises new questions: While the LIC model combining these modules achieves comparable overall performance, do they perform equally at each stage?
+If not, could an optimal configuration maximize their strengths at specific stages, further improving LIC performance without sacrificing efficiency?
+Through experiments, we identified the best configuration under the current framework, with detailed results in Sec.
+4.4.
+On such basis, we introduce the S2C-Hybrid model, an enhanced version of the S2CFormer-based LIC model.
+This hybrid model employs different S2C instances for different transform stages.
+Specifically, we employ convolution in the first stage when feature resolution is high, and apply attention mechanisms in the later two stages.
+This strategy improves R-D performance without increasing time complexity or slowing down coding speed.
+S2C-Hybrid offers an even more excellent trade-off between R-D performance and decoding speed.
+Previous SOTA Methods S2C Basic Instances S2C-Hybrid and Scaling-Up
+ELIC MambaVC FTIC MLIC++ TCM S2C-Identity S2C-Conv S2C-Attention Hybrid-S Hybrid-M Hybrid-L
+BD-rate [Kodak] -3.10 -8.11 -12.94 -11.97 -10.04 -9.70 -12.65 -12.56 -13.35 -13.85 -14.28
+BD-rate [CLIC] -0.84 - -10.21 -12.08 -8.60 -7.55 -10.98 -10.28 -12.07 -12.52 -12.88
+BD-rate [Tecnick] -7.41 - -13.89 -15.13 -10.42 -10.70 -14.48 -14.08 -15.21 -16.34 -17.20
+Params (M) 33.29 47.88 69.78 116.48 75.89 64.63 66.60 68.42 68.00 72.73 79.83
+FLOPs (T) 1.74 2.10 2.38 2.64 3.74 2.80 3.13 3.42 3.20 3.59 4.17
 […]
-Throughput (samples/s) $\uparrow$ 72.32 6.55 23.25 27.42 17.80 50.02 35.49 26.21 22.63
+Throughput(samples/s) 72.32 6.55 23.25 27.42 17.80 50.02 35.49 26.21 30.21 28.07 22.63
 Table 1: Comprehensive comparisons of LIC models.
@@
 S2C-Conv, -Attention, and -Hybrid models achieve superior R-D performance.
-R-D curves are provided in Fig.
-6-8.
 (b) Model Params.
-S2C-Conv and -Attention models have fewer parameters compared to previous SOTA methods.
+S2C-based models have fewer parameters compared to previous SOTA methods.
 (c) Efficiency.
-FLOPs assess computational complexity, but they do not always reflect the actual training and inference speed.
-While S2CFormer-based models do not exhibit many advantages in terms of FLOPs, they achieve over 30% faster decoding speed and significantly higher training throughput.
+S2C-based models achieve over 30% faster decoding speed and significantly higher training throughput.
+However, they are not superior in terms of FLOPs, and we will explain this issue in 4.5.
 4 Experiments
@@
 In the attention modules, the window size is set to 8, and depthwise convolution kernels are all of size 5.
-The latent representation $y$ consists of 320 channels, and the hyper-latent representation contains 192 channels.
-Additional details are provided in the supplementary material.
+We set {192, 192, 192, 320} for channel numbers $C_{1},C_{2},C_{3},C_{4}$.
 To enhance context mining and parameter aggregation within the entropy model of ELIC [21], we incorporate the S2C-Conv block and the S2C-Identity block, respectively.
@@
 We utilize the BD-rate [5] results to quantify the average bitrate savings.
-4.2 Rate-Distortion Performance
-In this section, we compare our models with VTM-21.0 [7, 6] and several leading LIC models, including ELIC [20], MambaVC [43], TCM [32], MLIC++ [22] and FTIC [27].
-The detailed results are presented in Table 1.
+We compare our models with VTM-21.0 [7, 6] and several leading LIC models, including ELIC [21], MambaVC [44], TCM [33], MLIC++ [23] and FTIC [28]. The detailed results are presented in Table 1.
+4.2 Basic S2CFormer Instances
 The Lower Bound of S2CFormers.
@@
 It surpasses MambaVC, and nearly matches the performance of TCM-large, as shown in Fig.
-4.
+5.
 This promising result suggests that MLP-based channel aggregation secures a solid baseline performance for LIC.
@@
 Although MLIC++ performs well at lower bit rates, it underperforms at higher bitrates, showing a noticeable gap compared to our models, as shown in Fig.
-4.
+5.
 This discrepancy arises from its focus on improving the entropy model, which benefits performance at lower bitrates.
@@
 The performances of our models suggest that complex spatial interaction design may be redundant and can be simplified.
-S2C-Conv and S2C-Attention show near-identical performance, indicating a specific choice of spatial interaction is not critical for LIC.
+S2C-Conv and S2C-Attention show near-identical performance, indicating a specific choice of spatial interaction is not that critical for LIC.
 Instead, FFN-based channel aggregation is the key factor that secures R-D performance.
@@
 Figure 9: Performance evaluation on the Tecnick dataset.
+4.3 Channel Aggregations
 Ablations of Channel Aggregation.
-In this section, we further analyze the critical role of Channel Aggregation within our S2CFormer architecture.
-To assess their impact, we first conducted an ablation study by removing all FFN modules from S2C-Conv and S2C-Attention.
-Fig.
-4 illustrates the R-D curves for this ablation study.
-Notably, pure separable convolution and window attention are significantly less effective than their FFN-enhanced counterparts.
-Specifically, adding FFNs to separable convolution improves BD-rate by -10.57%, while FFNs in S2C-Attention lead to an -11.14% improvement.
-These results highlight the substantial contribution of FFNs to the overall R-D performance of our LIC models.
-In contrast, S2C-Identity, which eliminates spatial interaction while retaining channel aggregation, approaches the performance of TCM-large, outperforming the two models without CA by -7.51% and -8.17% BD-rate respectively.
-This comparison underscores the greater importance of FFNs in LIC models.
-This ablation study strongly supports using FFNs for channel aggregation in future LIC models.
+In this section, we analyze the critical role of Channel Aggregation (CA) in the S2CFormer architecture.
+An ablation study was conducted by removing all FFN modules from S2C-Conv and S2C-Attention.
+Results in Fig.
+5 show that FFN-enhanced separable convolution and window attention significantly outperform their counterparts, improving BD-rate by -10.57% and -11.14%, respectively.
+This highlights the substantial impact of FFNs on R-D performance.
+In contrast, S2C-Identity, which retains CA but eliminates spatial interaction, approaches TCM-large’s performance, outperforming models without CA by -7.51% and -8.17% BD-rate.
+The study emphasizes the importance of FFNs in LIC models and supports their use for CA in future designs.
 Figure 10: Comparison of Different FFNs. Experiments on the Kodak dataset. We provide R-D curves for S2C-Conv and -Attention models with various FFNs for Channel Aggregation.
@@
 This figure presents visualizations of decompressed images from the Kodak dataset using various methods.
-Each subfigure is labeled with “Method |||| Bpp |||| PSNR |||| MS-SSIM”.
-We also examine the ERF [36] of analysis networks to demonstrate the significance of channel aggregation.
-Although it is commonly believed that the attention mechanism expands the receptive field and aggregates more non-local information, our findings challenge this view.
+Each subfigure is labeled with “Method |||| Bpp |||| PSNR |||| MS-SSIM".
+We analyze the Effective Receptive Field (ERF) [37] of networks to highlight the importance of channel aggregation.
+Contrary to the belief that attention mechanisms expand the receptive field and aggregate non-local information, our findings show that both S2C-Conv and S2C-Attention models exhibit smaller ERFs without channel aggregation.
 As shown in Fig.
-5, both the S2C-Conv and S2C-Attention models, when lacking channel aggregation, exhibit smaller ERF.
-In contrast, when channel aggregation is introduced, the ERF of both models expands to cover more areas of the input image.
-We also evaluate the Swin Transformer-based LIC model and observe that window shifting does not significantly enlarge the ERF.
-Similar to the S2C-based models, Swin-based models also rely on channel aggregation to broaden ERF.
-This demonstrates that channel aggregation is the key factor in activating more pixels in the original image and capturing additional information, which boosts the R-D performance of LIC models.
-These comparisons suggest that some complex and time-consuming spatial operations appear redundant.
+6, introducing channel aggregation significantly expands the ERF, covering more input image areas.
+Similarly, Swin Transformer-based LIC models rely on channel aggregation to broaden ERF, as window shifting alone does not significantly enlarge it.
+These results indicate that channel aggregation is crucial for activating more pixels and enhancing R-D performance, suggesting that certain complex spatial operations may be redundant.
 Effectiveness of Advanced Channel Aggregation.
@@
 We conduct comparison experiments on three proposed FFN structures—Vanilla-FFN, Additive-FFN, and Gated-FFN—within both the S2C-Conv and S2C-Attention models, as shown in Fig.
-9.
+10.
 For S2C-Conv, models utilizing Vanilla-FFN, Additive-FFN, and Gated-FFN outperform VTM-21.0 by -10.88%, -11.87%, and -12.65%, respectively.
@@
 These results demonstrate that the FFN structure remains a promising area for further investigation.
-Performance of S2C-Hybrid.
-We enhance model performance by integrating the advantages of different S2C-instantiations and advanced FFN.
-We increase $L_{2}$ and $L_{3}$ to 8 while keeping other configurations unchanged.
-To manage decoding latency and complexity, we utilize the S2C-Conv block in the initial stage to reduce complexity; in subsequent stages, we employ the S2C-Attention block.
-As shown in Fig.
-6-8, by employing only the fundamental S2CFormer blocks with simplified spatial interaction modules, our S2C-Hybrid model surpasses all other LIC models in R-D performance, exceeding VTM-21.0 by 14.28%, 12.88%, and 17.20% in BD-rate for the Kodak, CLIC, and Tecnick datasets, respectively.
-This result demonstrates the effectiveness of integrating various simple spatial interactions within a unified architecture to optimize both computational efficiency and model performance.
-Ablations on S2C-Hybrid For the three-stage main transform networks, we employ S2C-Conv for the first stage and S2C-Attention for the others, which is the best arrangement according to our experiments.
-As shown in Tab.2, we represent S2C-Conv and -Attention with “C” and “A” respectively.
-“[C,A,A]” appears to be the best, the reason for which may be that convolution can better capture local details at high resolution, while attention excels at global information when resolution drops.
+4.4 S2C-Hybrid
+Arrangement Ablations.
+Our experiment results in Tab.
+2 show that the optimal configuration for the three-stage main transform networks is to use S2C-Conv (denoted as “C”) in the first stage and S2C-Attention (denoted as “A”) in the subsequent stages.
+The “[C, A, A]” configuration performs best, likely because convolution layers excel at capturing local details at high resolutions, while attention mechanisms are more effective at extracting global information as resolution decreases.
+This highlights that while S2C-Conv and S2C-Attention exhibit similar rate-distortion performance, their distinct capabilities at different stages significantly impact overall performance when optimally arranged.
 Stages[1,2,3] [C,C,C] [A,A,A] [A,C,C] [C,C,A] [C,A,A]
@@
 Table 2: S2C-Hybrid Arrangement Ablations
-We also have experimented about the S2C-Hybrid model’s complexity in Tab.3.
-The base model Hybrid-S has 3 blocks for each stage.
-By stacking more blocks in the last two stages (adding blocks to the first stage greatly slows decoding), performance steadily improves.
-Stages[1,2,3] Hybrid-S[3,3,3] Hybrid-M[3,5,5] Hybrid-L[3,8,8]
-BD-rate (%) -13.35 -13.85 -14.28
-FLOPs (T) 3.20 3.59 4.17
-Decoding (s) $\downarrow$ 0.357 0.358 0.362
-Table 3: S2C-Hybrid Complexity Ablations
-Visual Comparison.  We present the visualization results in Fig. 10, comparing our method with the traditional codec VTM 21.0 and the advanced LIC model TCM.
+Performance and Scaling-Up.
+As shown in Table 1, the standard S2C-Hybrid model variant, Hybrid-S, which uses a configuration of {3, 3, 3} for stages $L_{1}$, $L_{2}$, and $L_{3}$, has achieved state-of-the-art performance across all three datasets while maintaining low decoding latency.
+Subsequently, we scaled up the S2C-Hybrid model to further explore the trade-off between model complexity and rate-distortion performance.
+Hybrid-M uses {3, 5, 5} blocks and Hybrid-L uses {3, 8, 8} blocks for stages $L_{1}$, $L_{2}$, and $L_{3}$.
+As reported in Table 1, stacking more blocks in the last two stages consistently improves performance.
+The RD curves for Hybrid-L are presented in Figures 79.
+Visual Comparison As illustrated in Fig. 11, comparisons with VTM-21.0 and TCM demonstrate that our proposed method achieves superior detail preservation.
 4.5 Model efficiency
-We calculate the FLOPs and decoding latency for various models using 2K images, and evaluate training throughput with a batch size of 8 and patch size of 256×256.
+We calculate the FLOPs and decoding latency for various models using 2K images, and evaluate training throughput with a batch size of 8 and patch size of 256$\times$256.
 Training throughput is directly correlated to the training time of LIC models.
@@
 FLOPs.
-For VAE-based LIC models, FLOPs are significantly influenced by the number of channels in the first stage.
-We directly set all channel numbers as 192, following the original setting of LIC model [2], as the optimal channel settings are beyond the scope of this study.
-Some leading methods have reduced the first-stage channel number to 128 or fewer [43, 27], which significantly lowers FLOPs.
-However, these models are still much slower than our S2CFormer-based models in both training and decoding speed, which are the primary concerns for LIC.
-5 Conclusion and Future Work
-We revisit the key factors that enhance the R-D performance of transformer-based LIC models.
-Specifically, we identify certain spatial operations in previous leading methods as redundant.
-To address this, we introduce S2CFormer, a novel paradigm for nonlinear transform blocks in LIC.
-This architecture consists of two core components: the Spatial Interaction module and the Channel Aggregation module.
-By shifting the focus of LIC from spatial interaction to channel aggregation, this approach aims to optimize performance while also improving speed.
-Through extensive experiments and comparisons, we demonstrate that FFN-based channel aggregation is critical for improving LIC performance.
-Models based on S2CFormer offer faster training speed and lower decoding latencies compared to existing methods while maintaining competitive R-D performance.
-Moreover, we suggest that FFN-based channel aggregation remains a promising avenue for future research.
-By combining simple spatial interaction operators with the proposed Gated-FFN for channel aggregation, our models achieve state-of-the-art R-D performance on the Kodak, Tecnick, and CLIC Professional Validation datasets, while also maintaining efficiency.
-This approach breaks the suboptimal trade-off between R-D performance and decoding latency inherent in previous methods.
+For VAE-based LIC models, the number of channels in the first stage significantly influences the FLOPs.
+In our work, we set all channel numbers to 192, following the classical setting of LIC models [2, 42, 3, 10, 21], since optimizing channel settings falls beyond the scope of this study.
+Although some leading methods reduce the first-stage channel count to 128 or fewer [44, 28] to lower FLOPs, these models remain considerably slower than our S2CFormer-based models in both training and decoding.
+This paper focuses on rebalancing spatial interactions and channel aggregation for low decoding latency.
+In this context, S2CFormer serves as a paradigm that permits flexible channel adjustments to reduce FLOPs.
+To demonstrate this, we propose a variant named Hybrid-T, designed for low-FLOP operation.
+Specifically, we set the channel numbers $C_{1}$, $C_{2}$, and $C_{3}$ to 96, 192, and 256, and the block numbers $L_{1}$, $L_{2}$, and $L_{3}$ to 3, 5, and 8, respectively.
+Hybrid-T achieves a SOTA BD-rate of -13.32% on the Kodak dataset, while its FLOPs are only 2.63T—a value relatively low compared to previous SOTA methods.
+These results indicate that our S2CFormer is not a fixed design, and we can achieve even more efficient S2C models by adjusting the channel numbers to find the optimal configuration in the future.
+5 Conclusion
+This paper tackles the trade-offs between decoding latency and rate-distortion performance in transformer-based LIC models.
+We emphasize the critical role of channel aggregation and demonstrate through extensive experiments that channel operations—rather than some time-consuming spatial interactions—are key to achieving high-performance LIC models.
+Based on these insights, we propose the S2CFormer paradigm, which streamlines spatial operations while substantially enhancing channel aggregation.
+Building on these findings, we evaluate several S2CFormer instances.
+Experimental results from S2C-Conv and S2C-Attention confirm that combining simpler spatial interactions with effective channel aggregation not only delivers state-of-the-art compression performance but also significantly accelerates decoding speed.
+Furthermore, the advanced S2C-Hybrid model, which integrates the strengths of multiple S2C variants, establishes a new benchmark for LIC by outperforming existing models on standard datasets such as Kodak, Tecnick, and CLIC Professional Validation.
+Overall, this work sets a new standard for efficient, high-performance LIC and lays a robust foundation for future research on optimized channel aggregation strategies.